OpenStudy (hitaro9):

Let ak >=0. Let the series ak diverge. Prove that the series ak/(1+ak) diverges

OpenStudy (hitaro9):

There's nothing super simple like a comparison that can be made as far as I can see. How would one go about proving this sort of question? Do you get in to definitions of series divergence or?

OpenStudy (hitaro9):

\[\sum_{k=1}^{\infty}a _{k}\div(a _{k}+1)\]

OpenStudy (hitaro9):



OpenStudy (hitaro9):

Hey, last problem. Thanks for your help on the last one @IrishBoy123. I'm just not sure what the general strategy is here.

OpenStudy (irishboy123):

compare like terms in each series

ganeshie8 (ganeshie8):

I think you may simply use the limit test : \(\lim\limits_{k\to\infty} \dfrac{a_k}{1+a_k} \ne 0 \implies \sum\limits_{k=0}^{\infty}\dfrac{a_k}{1+a_k} \text{ diverges}\)

OpenStudy (hitaro9):

Huh. How do we know ak / 1+ak doesn't go to zero? Couldn't ak go to zero and not converge (ie, 1/k)? Or do I have something off here?

OpenStudy (irishboy123):

yeah, tad confused here too we do not know that \(\lim\limits_{k\to\infty} a_k \ne 0\) as \(a_k\) might have failed that test and still diverged, ie it is poss that \(\lim\limits_{k\to\infty} a_k = 0\) plus \(\dfrac{\dfrac{a_k}{1+a_k}}{a_k} = \dfrac{1}{1 + a_k}\) and \(a_k >=0\) so the comparison test draws a blank

OpenStudy (anonymous):

$$ \sum \frac{\frac 1 n }{1 + \frac 1 n }= \sum \frac {1}{n+1 }= \infty $$

ganeshie8 (ganeshie8):

If \(\lim\limits_{k\to\infty} a_k = 0\), then limit comparison test does the job : \(\lim\limits_{k\to\infty} \dfrac{\dfrac{a_k}{1+a_k}}{a_k} = 1\) therefore both the series, \(\sum a_k \) and \(\sum\dfrac{a_k}{1+a_k}\) will have same fate.

ganeshie8 (ganeshie8):

looks im just repeating what irishboy has said

OpenStudy (irishboy123):

no, i chickened out of your second conclusion. can't remember why now :-)

OpenStudy (hitaro9):

So why does 1/1+ak -> that it diverges exactly? Sorry if this is obvious.

ganeshie8 (ganeshie8):

hmm question is not so clear... could you elaborate a bit

OpenStudy (anonymous):

limit comparison test: http://prntscr.com/8j65pv

OpenStudy (hitaro9):

Oh wait right. Thank you Jayz. Okay then that makes sense.

OpenStudy (hitaro9):

So basically we have two cases, one for which ak -> 0 and one for otherwise? If I understand correctly?

OpenStudy (anonymous):

:) no problem. ganeshie and irishboy did the hard work

OpenStudy (anonymous):

oh you mean if a_k converged, then yes

OpenStudy (hitaro9):

Okay yeah, my question was poorly phrased. I think I understand now. I'll go through and make it more wordy with respects to the limit test

OpenStudy (anonymous):

We get two statements for free using limit comparison. 1. Let ak >=0. Let the series ak diverge. Then the series ak/(1+ak) diverges 2. 1. Let ak >=0. Let the series ak converge. Then the series ak/(1+ak) converges

OpenStudy (hitaro9):

Thanks all of you for the help. I appreciate it.

OpenStudy (hitaro9):

Alright, I gotta get to class, thanks again for all your help. You're all amazing people.

OpenStudy (anonymous):

Just to go over the argument one more time to make this proof airtight. We are given ak>=0 and ∑ak diverges. Show that ∑ak / (1 + ak) diverges. proof: there are two cases lim ak ≠0 or lim ak = 0 1. If lim ak ≠0 , then since ak>=0 we have lim ak/(1 + ak) ≠0 and thus ∑ak / (1 + ak) diverges by divergence test. 2. If lim ak = 0 then lim [ ak / (1+ ak) ] / ak = lim ( 1 / (1 + ak) = 1 / (1 + 0) = 1 . by limit comparison test ∑ak / (1 + ak) diverges since ∑ak diverges